OpenStudy (anonymous):

I need to finish this problem and I'm not sure if I look for the lcm or the gcm. This is what I have so far: 4x(x-9)-5x(x-8)=-5 4x^2-36x-5x^2+40x+5=0

OpenStudy (anonymous):

4x(x-9)-5x(x-8) = 0 that distributed out is equal to 4x^2 - 36x - 5x^2 + 40x = 0 combine like terms -x^2 + 4x = 0 factor again -x(x - 4) = 0 and solve each factor for 0 -x = 0 x - 4 = 0 x = 0, 4

OpenStudy (anonymous):

When you combine like terms, are you adding or multiplying them? I don't remember the process.

OpenStudy (anonymous):

4x^2 - 36x - 5x^2 + 40x = 0 if you were to combine these terms together, it would look like this (basically a re-arranging of the equation): 4x^2 - 36x - 5x^2 + 40x = 0 (4x^2 - 5x^2) - (36x + 40x) = 0 (4-5)*(x^2 - (36-40)x = 0 -1x^2 - 4x = 0 -x^2 - 4x = 0 etc.

OpenStudy (anonymous):

and that should actually say this: (4-5)*x^2 - (36+40)x = 0 -1x^2 - 4x = 0 -x^2 - 4x = 0 from the 3rd line down. cause 40 was never negative! woops.

OpenStudy (anonymous):

Okay one more question then, -(36+40) would become a positve 4, or do you switch the signs inside of the paranthesis because there is a negative in front of them?

OpenStudy (anonymous):

I know I'm missing something simple here :(

OpenStudy (anonymous):

my error again... 4x^2 - 36x - 5x^2 + 40x = 0 <--- wrong (4x^2 - 5x^2) - (36x + 40x) = 0 <---wrong is actually 4x^2 - 36x - 5x^2 + 40x = 0 (4x^2 - 5x^2) - (-36x + 40x) = 0 <--- correct (-36 + 40 = 4 (-1)(x^2) + (-4)(x) = 0 -x^2 - 4x = 0 etc. so IN ALL. 4x^2 - 36x - 5x^2 + 40x = 0 (4-5)*x^2 - (-36+40)x = 0 (-1)x^2 - (4)x = 0 -x^2 - 4x = 0 -x^2 - 4x = 0 etc. man took me awhile to find that error. I didn't keep -36 negative, and then didn't multiply it by a negative factor again.

OpenStudy (anonymous):

O.o Lol, okay so when I simplify that I will have x= X^2 and x = 4x ?